OpenStudy (anonymous):

can someone please HELP me with this question? Graph the equation 4x+5y=20 Please show me step by step because in finding it hard to understand how to do these

OpenStudy (anonymous):

Solve the equation for y, which will give us the form y=mx+b 4x-4x+5y = 20-4x 5y=20-4x 5y/5 = 20/5 - 4x/5 y = 4 - 4x/5

OpenStudy (anonymous):

y = mx + b m = slope, b = y-intercept (where the line crosses the y-axis) The line will cross the y-axis at 4 or (0,4)

OpenStudy (anonymous):

From there use the slope, which is the number by x...-4/5. Start at the point you plotted (0,4)...go up 4 and to the left 5. You can also start at (0,4) and go down 4 and to the right 5.

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

YES thank you!